# NeuralOM: Neural Ocean Model for Subseasonal-to-Seasonal Simulation

### v2

@@ Appendix C Model Training
For the ocean-specific multi-stage framework, we apply Q sub-models to progressively capture subtle changes in ocean variables. We train the first sub-model with M steps finetune. For other sub-models, we train them with N steps finetune. In our experiment, we set Q=2, M=6, and N=10 as an example. Specifically, we train the first sub-model, denoted as $\mathcal{F}^{1}_{1}$ with 200 epochs using 1-step supervision. The learning rate is set to 1e-3. And we use the cosine annealing scheduler to adjust the learning rate until the model converges. After that, we finetune $\mathcal{F}^{1}_{1}$ with 10 epochs using 2-step supervision to get $\mathcal{F}^{2}_{1}$. Then, we finetune $\mathcal{F}^{2}_{1}$ with 10 epochs using 3-step supervision to get $\mathcal{F}^{3}_{1}$. The learning rate is set to 1e-6. And we use the cosine annealing scheduler to adjust the learning rate until the model converges. The model $\mathcal{F}^{4}_{1}$, $\mathcal{F}^{5}_{1}$, and $\mathcal{F}^{M}_{1}$ are finetuned same as $\mathcal{F}^{3}_{1}$. Then we freeze the parameters of $\mathcal{F}^{M}_{1}$ and train the second sub-model $\mathcal{F}^{1}_{2}$ with 200 epochs. The learning rate is set to 1e-6. And we use the cosine annealing scheduler to adjust the learning rate until the model converges. In practice, the model converges after a few epochs, and we stop the training early. The the $\mathcal{F}^{N}_{2}$ can be get step by step with same training produce as $\mathcal{F}^{1}_{2}$. Furthermore, the Q-th model $\mathcal{F}^{N}_{Q}$ can be trained step by step if a higher accuarcy is required. We train all baseline models and our NeuralOM using the same training framework. We will open-source all of our materials—including all training codes, detailed training procedures, testing codes, pre-trained weights, training logs, and more.
Appendix D Additional Results
@@
Additional quantitative results (RMSE and ACC across depth levels 109.73 m, 222.48 m, and 453.94 m) are shown in Figure 5. The RMSE and ACC represent the average performance for 240 ICs, starting from Jan. 1, 2020, and the interval of each IC is 1 day. Additional visual results are presented in Figure 6, Figure 7, Figure 8, Figure 9, Figure 10, and Figure 11, covering 17 important ocean variables across simulation time from 10 to 60 days, with the initial condition at Jun. 1, 2020. Specifically, for variables S, T, and SSH, where variations are subtle relative to their baseline values, anomalies are presented to highlight these minor but significant changes. Conversely, the original data are directly displayed for U and V. For clarity, examples of variable naming conventions are as follows: SSSa denotes sea surface salinity anomaly, Sa109 denotes salinity anomaly at 109.73 m, Sa222 denotes salinity anomaly at 222.48 m, and Sa453 denotes salinity anomaly at 453.94 m. U0 denotes the sea surface stream zonal velocity, U109 denotes the sea stream zonal velocity at 109.73 m, U222 denotes the sea stream zonal velocity at 222.48 m, and U453 denotes the sea stream zonal velocity at 453.94 m. Collectively, these supplementary results further demonstrate the efficacy of the proposed NeuralOM.

### v3

@@ C.2    Model Training
For the progressive residual correction framework, we apply Q sub-models to progressively capture subtle changes in ocean variables. We train the first sub-model with M steps finetune. For other sub-models, we train them with N steps finetune. In our experiment, we set Q=2, M=6, and N=10 as an example. Specifically, we train the first sub-model, denoted as $\mathcal{F}^{1}_{1}$ with 200 epochs using 1-step supervision. The learning rate is set to 1e-3. And we use the cosine annealing scheduler to adjust the learning rate until the model converges. After that, we finetune $\mathcal{F}^{1}_{1}$ with 10 epochs using 2-step supervision to get $\mathcal{F}^{2}_{1}$. Then, we finetune $\mathcal{F}^{2}_{1}$ with 10 epochs using 3-step supervision to get $\mathcal{F}^{3}_{1}$. The learning rate is set to 1e-6. And we use the cosine annealing scheduler to adjust the learning rate until the model converges. The model $\mathcal{F}^{4}_{1}$, $\mathcal{F}^{5}_{1}$, and $\mathcal{F}^{M}_{1}$ are finetuned same as $\mathcal{F}^{3}_{1}$. Then we freeze the parameters of $\mathcal{F}^{M}_{1}$ and train the second sub-model $\mathcal{F}^{1}_{2}$ with 200 epochs. The learning rate is set to 1e-6. And we use the cosine annealing scheduler to adjust the learning rate until the model converges. In practice, the model converges after a few epochs, and we stop the training early. The the $\mathcal{F}^{N}_{2}$ can be get step by step with same training produce as $\mathcal{F}^{1}_{2}$. Furthermore, the Q-th model $\mathcal{F}^{N}_{Q}$ can be trained step by step if a higher accuarcy is required. We train all baseline models and our NeuralOM using the same training framework. We will open-source all of our materials—including all training codes, detailed training procedures, testing codes, pre-trained weights, training logs, and more.
D    Additional Results
@@
Additional quantitative results (RMSE and ACC across depth levels 109.73 m, 222.48 m, and 453.94 m) are shown in Figure 6. Due to the relatively slow vertical variations in the ocean, the differences between adjacent layers at similar depths are relatively small. Therefore, selecting representative depth levels allows for a more intuitive assessment of the model’s ability to simulate the three-dimensional structure of the ocean. RMSE and ACC represent the average performance for 240 ICs, starting from Jan. 1, 2020, and the interval of each IC is 1 day. Additional visual results are presented in Figure 7, Figure 8, Figure 9, Figure 10, Figure 11, and Figure 12, covering 17 important ocean variables across simulation time from 10 to 60 days, with the initial condition at Jun. 1, 2020. Specifically, for variables S, T, and SSH, where variations are subtle relative to their baseline values, anomalies are presented to highlight these minor but significant changes. Conversely, the original data are directly displayed for U and V. For clarity, examples of variable naming conventions are as follows: SSSa denotes sea surface salinity anomaly, Sa109 denotes salinity anomaly at 109.73 m, Sa222 denotes salinity anomaly at 222.48 m, and Sa453 denotes salinity anomaly at 453.94 m. U0 denotes the sea surface stream zonal velocity, U109 denotes the sea stream zonal velocity at 109.73 m, U222 denotes the sea stream zonal velocity at 222.48 m, and U453 denotes the sea stream zonal velocity at 453.94 m. Collectively, these additional results further demonstrate the efficacy of the proposed NeuralOM.